OpenStudy (anonymous):

Simplify the expressions: (5 – 5i) – (–6 – 3i) a. –1 – 8i b. 11 – 2i c. –11 + 2i d. 9i (–2 – 2i)(–4 + 6i) a. 8 – 4i b. 8 – 12i c. 20 – 4i d. –4 – 4i

OpenStudy (yoloshroom):

So it's basically 5-5i +6+3i so 11-2i for the first one and use foil for the 2nd one

OpenStudy (yoloshroom):

and the second is 20-4i

OpenStudy (anonymous):

Okay, thank you so much. Have a wonderful night. c:

OpenStudy (dn_20):

So the first one: (5 – 5i) – (–6 – 3i), = 5 -5i +6 +3i = 11 -2i B is correct The second one (–2 – 2i)(–4 + 6i) FOIL = 8 -12i +8i -12i^2 =12 -4i +6 = 20 - 4i C is correct

OpenStudy (anonymous):

Ahh, I see thanks for explaining it. @DN_20